OpenStudy (anonymous):

Carlos and his friends decide to take a cab to the mall. The cab ride comes to $12.75. Write an equation that will help Carlos determine how much each friend pays. Let's let y=number of friends in the cab z=cost for each friend

OpenStudy (31356):

12.75/y = z make sure you note that Carlos is included in the"number of friends". Otherwise, it will be 12.75/(y +1) = z.

OpenStudy (31356):

Medal?

OpenStudy (31356):

Or: z = 12.75 / y Because, the total cost divided by the number of people present will tell you how much each person should pay. Here, the total cost is 12.75 and we will divide by number of people (y) to get the amount each person pays (z)

OpenStudy (31356):

Thanks!

OpenStudy (anonymous):

so the equation would look like \[z=\frac{ 12.75 }{ y }\]

OpenStudy (31356):

Yes

OpenStudy (anonymous):

The shampoo you like to use is on sale for $3.69 per bottle.Write an equation that will help you determine how much you hsave by purchasing the bottle of shampoo on sale. Let's let p=regular price c=cost savings